# FCS-HGNN: Flexible Multi-type Community Search in Heterogeneous Information Networks

### v2

@@ 6.3. Ablation Study and Parameter Sensitivity Analysis
Refer to caption (a) DBLP Refer to caption (b) ACM
Refer to caption (a) DBLP Refer to caption (b) ACM
Refer to caption (a) DBLP Refer to caption (b) ACM
Figure 6. Ablation Study of EA.
Figure 7. Ablation Study of Query Enc.
@@
Refer to caption (a) DBLP Refer to caption (b) ACM
Refer to caption (a) DBLP Refer to caption (b) ACM
Refer to caption (a) DBLP Refer to caption (b) ACM
Figure 9. Parameter Analysis of $d_{max}$.
Figure 10. Parameter Analysis of $f_{0}$.
@@
Then, we investigate the impact of the hyperparameter fanouts $f_{0}$ and $f_{1}$ in Algorithm 3. For the analyses, $f_{0}$ is varied within {5,10,15,20} with a controlled $f_{1}=10$, and vice versa, where $f_{1}$ is varied within the same range while maintaining $f_{0}=20$. Figure 11 and 11 illustrate the performance with varying $f_{0}$ or $f_{1}$. We observe that the model is not sensitive to the fanouts $f_{0}$ or $f_{1}$, which is consistent with our previous analysis. Due to the random sampling of neighbors in each epoch, the model is able to learn node features and community structure in an efficient and stable manner.
@@ 7. CONCLUSIONS
In this paper, we propose a novel query-driven heterogeneous graph neural network, named FCS-HGNN, for multi-type community search in HINs. Compared to existing methods, it can flexibly identify both single-type and multi-type communities without relying on predefined meta-paths or user-specific relationship constraints, significantly alleviating the burden on users. FCS-HGNN dynamically consider the contribution of each relation instead of treating them equally, thus capturing more fine-grained heterogeneous information. Furthermore, to improve efficiency on large-scale graphs, we introduce a neighbor sampling algorithm to enhance training efficiency and propose the depth-based heuristic search strategy to improve query efficiency. We thoroughly evaluate our proposed methods on five real-world HINs and demonstrate its significant superiority in terms of effectiveness and efficiency.
@@ Acknowledgements.
We would like to express our gratitude to the teams led by Lei Chen (Jian et al., 2020) and Yixiang Fang (Jiang et al., 2022a) for laying the groundwork with their outstanding research achievements in this field. Fangda Guo is the corresponding author.

### v3

@@ 5.3. Ablation Study
(a) DBLP
(b) ACM
@@
(a) DBLP
(b) ACM
@@
Then, we investigate the impact of the hyperparameter fanouts $f_{0}$ and $f_{1}$ in Algorithm 2. For the analyses, $f_{0}$ is varied within {5,10,15,20} with a controlled $f_{1}=10$, and vice versa, where $f_{1}$ is varied within the same range while maintaining $f_{0}=20$. Figure 9 and 9 illustrate the performance with varying $f_{0}$ or $f_{1}$. We observe that the model is not sensitive to the fanouts $f_{0}$ or $f_{1}$, which is consistent with our previous analysis. Due to the random sampling of neighbors in each epoch, the model is able to learn node features and community structure in an efficient and stable manner.
@@ 6. CONCLUSIONS
In this paper, we propose a novel query-driven heterogeneous graph neural network, named FCS-HGNN, for multi-type community search in HINs. Compared to existing methods, it can flexibly identify both single-type and multi-type communities without relying on predefined meta-paths or user-specific relationship constraints, significantly alleviating the burden on users. Moreover, FCS-HGNN dynamically consider the contribution of each relation instead of treating them equally, thus capturing more fine-grained heterogeneous information. Furthermore, to improve efficiency on large-scale graphs, we introduce a neighbor sampling algorithm to enhance training efficiency and propose the depth-based heuristic search strategy to improve query efficiency. We thoroughly evaluate our proposed methods on five real-world HINs and demonstrate its significant superiority in terms of effectiveness and efficiency.
@@ Acknowledgements.
This work was supported by the National Key R&D Program of China under grant number 2023YFC3305303, the National Natural Science Foundation of China under grant number 62372434, 62302485, China Postdoctoral Science Foundation (No. 2022M713206) and CAS Special Research Assistant Program.